### v1

# ICon: In-Context Contribution for Automatic Data Selection

Abstract
Data selection for instruction tuning is essential for improving the performance of Large Language Models (LLMs) and reducing training cost. However, existing automated selection methods either depend on computationally expensive gradient-based measures or manually designed heuristics, which may fail to fully exploit the intrinsic attributes of data. In this paper, we propose In-context Learning for Contribution Measurement (ICon), a novel gradient-free method that takes advantage of the implicit fine-tuning nature of in-context learning (ICL) to measure sample contribution without gradient computation or manual indicators engineering. ICon offers a computationally efficient alternative to gradient-based methods and reduces human inductive bias inherent in heuristic-based approaches. ICon comprises three components and identifies high-contribution data by assessing performance shifts under implicit learning through ICL. Extensive experiments on three LLMs across 12 benchmarks and 5 pairwise evaluation sets demonstrate the effectiveness of ICon. Remarkably, on LLaMA3.1-8B, models trained on 15% of ICon-selected data outperform full datasets by 5.42% points and exceed the best performance of widely used selection methods by 2.06% points. We further analyze high-contribution samples selected by ICon, which show both diverse tasks and appropriate difficulty levels, rather than just the hardest ones. [^1] Project page: https://annayang2020.github.io/ICon_Data_Selection/
ICon: In-Context Contribution for Automatic Data Selection
[…]
1 Introduction
Figure 1: Overview of the ICon Method. The diagram illustrates the two key components of the method: Contribution Quantification with the ICon Score and ICon-guided Selection Paradigm Training. This approach enables gradient-free, bias-reduced data selection with scalable inference.
Recent advancements in large language models (LLMs), such as GPT-4o Hurst et al. (2024), have demonstrated strong capabilities in both understanding and generation. These models are capable of performing a wide range of tasks Liu et al. (2023b); Dong et al. (2023); Team et al. (2024); Guo et al. (2025), often showing impressive flexibility and adaptability. Instruction tuning Mishra et al. (2021); Wei et al. (2021); Sanh et al. (2022) has emerged as a powerful approach to enhance the performance of such models by finetuning them to better follow human instructions. While traditional instruction tuning relies on amassing vast datasets Köpf et al. (2023); Chung et al. (2024), recent study Zhou et al. (2023a) manually selecting data has shown that training with carefully selected subsets of high-quality data can achieve superior performance with reduced computational cost.
Automated data selection methods are essential, as manually selecting the most valuable training data is time-consuming and impractical. Previous studies on automated data selection Xia et al. (2024); Zhang et al. (2024) have explored gradient-based approaches to estimate sample value, but these methods are often computationally expensive and inefficient. Other works propose lightweight alternatives based on manually designed textual features Liu et al. (2023c); Bukharin and Zhao (2023); Zhang et al. (2025) or simple heuristics such as difficulty Li et al. (2023a, 2024b) and uncertainty Kung et al. (2023). However, these methods fail to directly measure the training impact of samples and introduce human inductive bias into the selection process. Specifically, manually defined indicators, such as difficulty, uncertainty, or topic coverage, can not accurately capture the true contribution of a sample, and reliance on predefined metrics limits the ability to fully exploit the potential of the training data.
In this paper, we introduce In-Context Learning for Contribution Measurement (ICon), a novel selection method that captures the contribution of individual samples while avoiding human inductive bias and costly gradient computation. Inspired by in-context learning (ICL), our method takes advantadge of recent findings Dong et al. (2022); Dai et al. (2023); Zheng et al. (2023); Yin et al. (2024) that ICL simulates parameter updates via the attention mechanism, effectively acting as implicit fine-tuning without the need for explicit gradient updates. Our key intuition is that this implicit finetuning nature allows us to directly assess the contribution of individual samples, without the need for manually defined indictor engineering. Additionally, since ICon does not require explicit training or gradient computation, it enables highly efficient selection, significantly reducing computational overhead compared to traditional gradient-based methods.
Our approach leverages ICL to measure sample contribution through three components: 1) an assessment set to provide a reliable performance reference, 2) ICon scores that quantify contribution via ICL’s implicit fine-tuning, and 3) a lightweight selection paradigm for scalable data selection. The ICon score captures each sample’s contribution without costly gradient computation or human-designed heuristics. A higher positive score indicates a greater contribution, while a lower negative score reflects more detrimental effects. Finally, the ICon-guided selection paradigm is trained efficiently with LoRA Hu et al. (2021) based on the target model, reducing the selection complexity to linear inference calls.
We demonstrate the effectiveness of ICon across multiple models and evaluation settings. Experiments on LLaMA3.1-8B Grattafiori et al. (2024), Qwen2.5-3B Yang et al. (2024), and LLaMA2-7B Lyu et al. (2024) are conducted using 12 widely used benchmarks and 5 pairwise comparison test sets. Models trained on a small fraction of ICon-selected data outperform their full-data counterparts. Notably, with only 15% of the data, the LLaMA3.1-8B and Qwen2.5-3B models improve by 5.42 and 1.24 percentage points, respectively; LLaMA2-7B achieves a 0.65-point gain using just 5%. ICon also surpasses widely used selection methods. For example, ICon improves LLaMA3.1-8B’s average benchmark score by 2.06 points over the best prior method. Furthermore, we also analyze the impact of data scale, cross-dataset generalization, and the properties of high-contribution samples selected by ICon. The main contributions of our paper are as follows:
  • We propose ICon, a novel data selection method inspired by ICL. By leveraging its implicit fine-tuning nature, ICon enables gradient-free and bias-reduced measurement of each sample’s contribution to model performance.
  • Our approach avoids gradient updates and auxiliary models, reducing computational cost and design complexity. Once trained with LoRA, the ICon-guided selection paradigm enables data selection through linear-complexity inference calls, making it highly scalable.
  • We demonstrate the effectiveness of ICon through experiments on multiple models, 12 widely used benchmarks, and 5 pairwise test sets, showing consistent performance gains over baselines.
2 Related Work
Data Selection
Data selection aims to construct optimal datasets to improve model performance Murphy (2012), reduce training cost Suárez et al. (2019); Sorscher et al. (2022), mitigate undesirable model behaviors Longpre et al. (2024), and ensure evaluation quality Oren et al. (2024). With the rise of LLMs, it plays a central role across various training stages, including pretraining, instruction tuning, alignment, in-context learning, and task-specific finetuning. Pretraining selection focuses on filtering large-scale raw data Soldaini et al. (2024); Penedo et al. (2023); Instruction tuning often uses manual Zhou et al. (2023a) or automated methods, with the latter detailed in the next section; alignment involves manual curation Bai et al. (2022), model-based evaluation Wang et al. (2023b); Cui et al. (2023), and reward model re-weighting Touvron et al. (2023); Pace et al. (2024); in-context learning emphasizes demonstration choice Xu and Zhang (2024); Luo et al. (2024) and order Lu et al. (2021); and task-specific tuning leverages utility-based Ivison et al. (2023) or empirical methods Grangier and Iter (2022).
Data Selection for Instruction Tuning
Instruction tuning Wei et al. (2021); Liu et al. (2023a); Xu et al. (2024) is a common approach to enhance instruction-following capabilities. Recent work Zhou et al. (2023a) shows that carefully selecting a smaller, higher quality dataset can improve both the effectiveness and efficiency of instruction tuning. Existing automatic data selection methods fall into three main categories: (1) selection based on textual features like helpfulness, quality, or complexity Zhang et al. (2025); Cao et al. (2023); Chen et al. (2023); Liu et al. (2023c); Bukharin and Zhao (2023); Minghao et al. (2024); (2) selection using training-time gradients to capture model-specific characteristics Xia et al. (2024); Zhang et al. (2024); Pan et al. (2024); Chang et al. (2024); (3) selection guided by feedback during inference, which is model-aware and much more efficient Du et al. (2023); Kung et al. (2023); Li et al. (2023a, 2024b); Liu et al. (2024); Wang et al. (2024); Li et al. (2024a). The first and third approaches rely on handcrafted heuristics for scoring and filtering, introducing human inductive bias, while the second demands substantial computational cost. ICon addresses the limitations of existing approaches by avoiding both costly gradient computations and manually designed heuristics. It leverages ICL to directly measure sample contribution in a model-aware, efficient manner.
3 Methodology
To efficiently assess the contribution of each sample, we draw on recent advances in in-context learning (ICL), which has been recognized as a form of implicit finetuning without explicit gradient updates Dai et al. (2023); Yin et al. (2024). Specifically, ICL leverages the model’s attention mechanism to simulate parameter updates by processing demonstration examples, thereby adapting the model’s behavior within the given context. The key intuition is that in-context learning can effectively simulate the training impact of individual samples, enabling an assessment of their influence on model performance without the high computational cost of finetuning.
Building on this insight, we propose the ICon method, an efficient data selection method that harnesses ICL’s implicit finetuning nature to assess each sample’s contribution to overall model performance. As shown in Figure 1, ICon comprises three phases: (1) constructing a representative assessment set, (2) computing the proposed ICon score to quantify sample contribution, and (3) learning an ICon-guided selection paradigm to curate data for final instruction tuning.
3.1 Assessment Set Construction
The goal of this phase is to construct an assessment set that reliably reflects the model’s overall capabilities. To ensure diversity and reduce bias, we sample data from three sources: ChatGPT-generated, GPT-4-generated, and human-authored data, resulting in a high-quality set covering a range of task types and instruction complexities. The assessment set is used solely for assessing model performance, with no overlap with training data or downstream benchmarks, ensuring independent evaluation and fair comparison. Formally, we define the assessment set as $D_{a}$, consisting of $n$ instruction tuning task samples, where each is represented as a pair $(x^{a},y^{a})$. Here, $x=\text{map}(\text{Instruction},[\text{Input}])$ represents the full instruction, which is formed by concatenating the instruction and optional input, and $y$ denotes the corresponding response. Similarly, we also define the candidate training set as $D_{t}=\{(x^{t}_{1},y^{t}_{1}),(x^{t}_{2},y^{t}_{2}),\ldots,(x^{t}_{m},y^{t}_{m})\}$, where $m$ is the number of candidate training instruction-response pairs.
3.2 Contribution Quantification with ICon Score
In this phase, we introduce the ICon score, a metric for estimating the contribution of individual samples to overall model performance. The core idea is to leverage ICL to evaluate contribution of sample. ICL enables models to adapt to new information presented within a prompt, providing an efficient approach to assess the contribution of individual samples on task performance. We adopt perplexity as the assessment metric, as it directly evaluates the likelihood of response tokens and serves as a smoother indicator of model performance compared to accuracy Li et al. (2023a, 2024b); Jiang et al. (2024). For a task sample $S_{i}=(x^{a}_{i},y^{a}_{i})\in D_{a}$, where $D_{a}$ is the assessment set, $y^{a}_{i}$ contains $N$ tokens and the $k$-th token is denoted as $y^{a}_{i,k}$, the model’s performance under parameters $\theta$ on task $i$ is defined as:
$\text{PPL}_{\theta}(S_{i})=\text{PPL}_{\theta}(y_{i}^{a}|x_{i}^{a})$ (1)
$=\exp\left(-\frac{1}{N}\sum_{k=1}^{N}\log p_{\theta}(y^{a}_{i,k}|x^{a}_{i},y^{a}_{i,1},...,y^{a}_{i,k-1})\right)$
Then, We select a sample $T_{j}=(x^{t}_{j},y^{t}_{j})$ from $D_{t}^{\prime}$, a subset of the training set $D_{t}$. This sample is inserted into the context during inference to assess the model’s performance on an assessment sample $S_{i}$ after implicitly learning from $T_{j}$. The model’s performance on $S_{i}$, given $T_{j}$ in context is denoted as $\text{PPL}_{\theta}(S_{i}|T_{j})=\text{PPL}_{\theta}(y^{a}_{i}\mid T_{j},x^{a}_{i})$. Since perplexity is affected by both content and input length, we introduce a fairness adjustment to neutralize length effects. Specifically, we define the baseline performance on $S_{i}$ without the influence of $T_{j}$ by using a randomly generated meaningless sequence $T_{j}^{\text{rand}}$, which matches the length of $T_{j}$. The model’s baseline performance on $T_{i}$ is then denoted as $\text{PPL}_{\theta}(S_{i}|T_{j}^{\text{rand}})=\text{PPL}_{\theta}(y^{a}_{i}\mid T_{j}^{\text{rand}},x^{a}_{i})$.
Using these values, we define the task-level ICon score of sample $S_{j}$ on task $i$ under parameters $\theta$ as shown in Equation 2, where $\epsilon$ is a small constant to avoid division by zero. A positive task-level ICon score indicates that $S_{j}$ contributes positively to model performance on task $i$, while a negative score suggests a detrimental effect. We then compute the global ICon score by aggregating the task-level scores across all assessment tasks, as shown in Equation 3. This global score allows us to rank samples based on their overall contribution. A higher positive global ICon score implies a stronger positive impact on model overall performance, while a lower negative score indicates the opposite.
[…]
Figure 2: Pairwise comparison results (win-tie-lose) between models trained on ICon-selected subsets and models trained on the full Alpaca dataset across five evaluation benchmarks: Vicuna, Koala, WizardLM, SInstruct, and LIMA. ICon-selected data enhances instruction-following ability with fewer samples, as evidenced by pairwise comparisons.
3.3 ICon-guided Selection Paradigm Training
While the ICon score effectively captures a sample’s contribution, it requires multiple inference calls per candidate training sample. To further improve efficiency, we train an ICon-guided selection paradigm using the computed global ICon scores. This allows us to reduce the selection complexity to $O(m)$ inference calls, where $m$ is the total number of candidate samples. Specifically, we use a subset of the training data and label the top $K\%$ of samples by global ICon score as high-contribution. A selection model is then trained using LoRA on the target model to classify high-value samples. This paradigm is applied to the full candidate set to select the most contributive samples for instruction tuning, thereby enhancing the model’s ability to follow instructions.
[…]
IFEval AlpacaEval GLUE GPQA MMLU TQA ARC BBH HS LQA MuSR WG (×$\times 10^{12}$)
LLaMA3.1-8B
FULL 15.68 18.66 54.48 27.27 43.26 39.60 40.53 36.46 51.42 26.73 37.51 63.77 37.95 40.12
[…]
ICon (5%) 16.21 28.96 56.91 27.27 53.70 40.51 47.18 44.23 58.44 24.42 37.75 69.46 42.09(+4.14) 3.06
ICon (10%) 16.08 27.99 57.86 26.76 53.57 42.63 45.48 42.81 56.27 26.57 42.01 66.93 42.08(+4.13) 5.62
ICon (15%) 21.16 29.89 60.58 30.81 52.30 43.16 44.80 44.15 55.56 28.42 41.64 67.96 43.37(+5.42) 8.37
Qwen2.5-3B
FULL 32.34 28.74 68.98 31.82 61.71 45.42 46.42 45.05 55.13 35.79 44.51 63.85 46.65 36.46
[…]
ICon (5%) 20.85 36.25 68.34 27.88 64.73 43.26 53.50 47.80 58.10 32.10 39.72 69.06 46.80(+0.15) 2.27
ICon (10%) 25.28 37.05 69.26 34.85 63.67 43.35 50.00 47.00 57.85 35.79 40.25 68.27 47.72(+1.07) 5.28
ICon (15%) 23.07 30.11 70.46 34.34 64.09 46.90 51.71 47.68 58.16 38.40 41.96 67.80 47.89(+1.24) 6.85
LLaMA2-7B
FULL 15.48 21.76 53.98 26.26 43.69 40.48 41.72 36.18 53.98 25.03 41.11 62.82 38.54 48.72
[…]
ICon (5%) 7.78 31.47 51.93 23.23 36.95 45.72 40.53 36.07 57.57 23.96 40.62 66.77 38.55(+0.01) 3.26
ICon (10%) 11.86 29.76 50.62 25.76 37.30 45.19 41.72 35.38 57.39 24.88 39.82 66.38 38.84(+0.30) 6.16
ICon (15%) 8.06 26.83 54.67 27.78 36.39 43.89 42.66 34.74 57.01 23.50 39.80 66.06 38.45(-0.09) 9.18
Table 1: Performance of LLaMA3.1-8B, Qwen2.5-3B, and LLaMA2-7B in terms of FLOPs and evaluation benchmarks for instruction following, knowledge, and reasoning. The datasets TQA, HS, LQA, and WG correspond to TruthfulQA, HellaSwag, LogicQA, and WinoGrande, respectively. Models trained on ICon-selected data outperform those trained on full datasets on evaluation benchmarks, achieving better performance with fewer samples.
4 Experiments Setup
4.1 Datasets
Training Dataset
In this paper, we use the classic Alpaca dataset Taori et al. (2023), one of the original instruction datasets created by Stanford University. The Alpaca dataset consists of 52,002 instruction-following samples, generated with TextDavinci-003 through the self-instruction approach Wang et al. (2022). The LLM trained on this dataset demonstrate a good instruction-following capability. However, its heavy dependence on TextDavinci-003 has raised concerns among researchers regarding the quality of instruction data. The WizardLM dataset Xu et al. (2023) utilizes the Evol-Instruct algorithm to enhance the quality and complexity of instruction data. We conduct experiments using WizardLM-70K to verify the generalization of our method across different datasets.
Assessment Dataset
We consider the impact of a training candidate sample on the model’s performance over the assessment dataset as an indicator of its overall influence on model capability, as constructed in Section 3.1. To reduce bias and ensure diversity, we sample data from three sources: OpenOrca-GPT3.5 Lian et al. (2023) containing ChatGPT-generated data, OpenOrca-GPT4 containing GPT-4-generated data, and Dolly-15K Mike et al. (2023) containing human-generated data. From these three datasets, we randomly select 1,020 instructions to construct the assessment dataset.
4.2 Implementation Details
For experiments on the LLaMA3.1-8B Grattafiori et al. (2024), Qwen2.5-3B Yang et al. (2024), and LLaMA2-7B Lyu et al. (2024) models, we all follow the original Alpaca [^2] https://github.com/tatsu-lab/stanford_alpaca training configuration using the Alpaca codebase. Models are trained for three epochs with the Adam optimizer Kingma and Ba (2014), a learning rate of $2\times 10^{-5}$, and a batch size of 128. The maximum input length is set to 512 for Alpaca dataset and 2048 for WizardLM dataset.
4.3 Evaluation Metrics
4.3.1 Benchmark Evaluation
To evaluate the effectiveness of ICon, we report performance on several widely used benchmarks. These benchmarks fall into three categories, assessing different capabilities: instruction following, knowledge, and reasoning. Instruction following evaluation includes two key benchmarks: IFEval Zhou et al. (2023c) and AlpacaEval Li et al. (2023b). AlpacaEval provides an LLM-based automatic evaluation on the AlpacaFarm set Dubois et al. (2023), where we use GPT-4 Achiam et al. (2023) as both the response generator and the evaluator comparing model outputs with reference responses. Knowledge evaluation includes four key benchmarks: GLUE Wang et al. (2018), GPQA Rein et al. (2024), MMLU Hendrycks et al. (2020), and TruthfulQA Lin et al. (2021). Reasoning evaluation includes six key benchmarks: ARC Clark et al. (2018), BBH Suzgun et al. (2022), HellaSwag Zellers et al. (2019), LogiQA Liu et al. (2020), MuSR Sprague et al. (2023), and Winogrande Sakaguchi et al. (2021).
4.3.2 Pairwise Comparison
The performance on several pairwise comparison test sets for instruction following is also provided, including WizardLM Xu et al. (2023), Self-Instruct Wang et al. (2022), Vicuna Chiang et al. (2023), Koala Vu et al. (2023), and LIMA Zhou et al. (2023b). These sets contain 218, 252, 80 and 300 human-curated instruction data, respectively, across domains such as math, coding, writing, and general knowledge. We use GPT-4 Hurst et al. (2024) as the judge. For each instruction in the test sets, candidate models generate responses, which are then scored by GPT-4 in a pairwise fashion. The judge rates each response from 1 to 10 based on factors such as relevance and accuracy. To mitigate positional bias Wang et al. (2023a), we submit each response pair twice with reversed order. A model is considered to win only if it does not lose in both orders, following these rules: Win: wins both, or wins one and ties the other; Tie: ties both, or wins one and loses one; Lose: loses both, or loses one and ties the other.
5 Experiment Results
5.1 Main Results
In this section, we present the evaluation benchmark results shown in Table 1. The models are trained using different proportions of Alpaca data selected by the ICon method, including 1%, 5%, 10%, and 15%, corresponding to 520, 2,600, 5,200, and 7,800 samples, respectively. "All" denotes models trained on the full, unfiltered dataset. Experiments were conducted on multiple models, including LLaMA3.1-8B, Qwen2.5-3B, and LLaMA2-7B. Across all benchmarks, models trained on a small fraction of ICon-selected data often match or surpass the performance of those trained on the full dataset, and consistently outperform them in overall metrics. For example, the LLaMA3.1-8B model trained on just 15% of ICon-selected data exceeds the full-data baseline by 5.42 percentage points. Similarly, the Qwen2.5-3B model with 15% ICon-selected data and the LLaMA2-7B model with only 1% achieve overall gains of 1.24% and 0.65%, respectively. Notably, these ICon-trained models require only about 1/5 to 1/100 of the FLOPs compared to full-data training, demonstrating significantly lower resource consumption. These results underscore the effectiveness of ICon-guided selection in enhancing instruction tuning with substantially less data.
The results of the pairwise evaluation, including detailed win-tie-lose statistics on the Vicuna, Koala, WizardLM, SInstruct, and LIMA test sets, are presented in Figure 2. We focus on the best-performing models from the evaluation benchmarks: LLaMA3.1-8B, Qwen2.5-3B, and LLaMA2-7B, each trained on 7,800, 7,800, and 520 ICon-selected samples, respectively. All three models consistently outperform their counterparts trained on the full Alpaca dataset across all five test sets. These results further validate the effectiveness of ICon-guided selection in improving instruction-following ability, even with significantly less training data.
5.2 Comparison with Other Methods
Comparison Pairwise Winning Score $\uparrow$ Overall
Test Set Vicuna WizardLM LIMA SInstruct Koala
ICon vs. Random 1.2125 1.4541 1.3233 1.5159 1.3389 1.3922
ICon vs. Low PPL 1.3375 1.3945 1.4533 1.3690 1.3333 1.3903
ICon vs. Top PPL 1.9126 1.9125 1.8567 1.8452 1.7778 1.8556
ICon vs. Alpagasus Chen et al. (2023) 1.1250 1.1789 1.1033 1.3611 1.3389 1.2252
ICon vs. Deita Liu et al. (2023c) 1.3375 1.2661 1.2733 1.3730 1.3889 1.3214
ICon vs. Superfilter Li et al. (2024b) 1.0375 1.0780 1.0100 1.1627 1.0944 1.0786
Table 2: Comparison with other methods on Vicuna, WizardLM, LIMA, SInstruct, and Koala test set. The pairwise Winning Scores are calculated between models using our method and other methods. All the comparisons are performed by GPT-4, and the values that are greater than 1.0 represent our models are better and vice versa. The models are trained on LLaMA3.1-8B with 7,800 selected samples. ICon consistently outperforms widely used data selection methods on pairwise comparison test sets.
In this subsection, we compare our method with several widely used data selection baselines on the Alpaca dataset, as shown in Table 2 and Table 4. "Random" selects samples at random. "Low PPL" selects samples with the lowest perplexity, while "Top PPL" selects those with the highest perplexity. "Alpagasus" uses GPT-3.5-Turbo to score sample responses based on dimensions such as helpfulness and accuracy. "Deita" uses a set of models fine-tuned from LLaMA and Mistral to evaluate sample quality, complexity, and diversity. "Superfilter" selects samples using IFD scores that reflect instruction difficulty, computed using GPT-2. Superfilter assumes that IFD scores from smaller models are as reliable as those from larger ones.
[…]
6 Analysis
Optimal Data Scale for Instruction Tuning
We further investigate how the scale of selected data affects model performance. Using the LLaMA3.1-8B model as a case study, Figure 3 presents the results of training on ICon-selected subsets of the Alpaca dataset at varying proportions: 1%, 5%, 10%, 15%, 20%, 25%, 30%, 50%, 75%, and 100%. The average score across all evaluation benchmarks generally rises and then declines, peaking at 15%, demonstrating that a small, carefully selected subset can outperform models trained on much larger datasets. Based on this observation, we select the model trained on 15% (7,800 samples) of ICon-selected data as the representative model for subsequent comparisons. Detailed results for each selection scale are provided in Appendix B. These results show that adding more data beyond 15% does not lead to further improvements, highlighting the efficiency of ICon in identifying the most contributive samples.
Figure 3: Model performance (average score) on evaluation benchmarks with varying proportions of ICon-selected Alpaca data. The models are trained based on LLaMA3.1-8B. The dataset is Alpaca dataset. As data scale grows, performance generally improves and then declines, with the best result at 15%.
Generalization of ICon on different Datasets
IF KN RS Avg
FULL 33.61 45.46 44.37 42.94
[…]
ICon (5%) 43.21 48.28 48.36 47.48
ICon (10%) 36.89 48.15 47.18 45.79
ICon (15%) 35.24 46.29 45.76 44.18
Table 3: Evaluation benchmark results of models trained on WizardLM dataset and its ICon-selected subsets. The ‘IF’, ‘KN’, and ‘RS’ correspond to average scores on Instruction Following, Knowledge, and Reasoning, respectively. Results on WizardLM dataset confirm the generalization of ICon across datasets.
IF KN RS Avg
Low PPL 19.35 44.45 46.53 41.31
Top PPL 3.28 41.72 44.69 36.80
Alpagasus 16.39 45.14 45.29 40.42
Deita 15.20 45.46 45.92 40.65
Superfilter 20.25 44.76 45.45 41.02
ICon (Ours) 25.52 46.71 47.09 43.37
Table 4: Performance of different methods on evaluation benchmarks. The ‘IF’, ‘KN’, and ‘RS’ correspond to average scores on Instruction Following, Knowledge, and Reasoning, respectively. The models are trained on LLaMA3.1-8B with 7,800 selected samples. ICon outperforms widely used data selection methods on pairwise evaluation benchmarks.
In addition to the main experiments conducted on the Alpaca dataset, we also evaluated the generalization of the ICon method to different datasets. Specifically, we performed experiments on the WizardLM dataset. The results, shown in Table 6 and detailed in Appendix C, cover selection scales of 1%, 5%, 10%, 15%, and ALL, corresponding to 520, 2,600, 5,200, 7,800, and 70K samples. As shown, models trained on ICon-selected subsets consistently outperform those trained on the full dataset across the Instruction Following, Knowledge, and Reasoning tasks, achieving higher average scores and overall performance. The model trained on just 5% of the data achieves the best results, with a 4.54 percentage point improvement in average evaluation benchmark score. These findings demonstrate the effectiveness of ICon across datasets. Notably, the selection paradigm, trained on the Alpaca subset, is reused here without modification, demonstrating its robust generalization.
Characteristics of High-Contribution ICon Samples
Figure 4: Visualization using t-SNE on sample embeddings from the Alpaca dataset. Red points represent samples with the top 15% ICon high-contribution scores and gray points represent other samples from the dataset. High-contribution ICon samples exhibit diverse characteristics.
Figure 5: Difficulty distribution of top 15% high-contribution samples vs. full Alpaca dataset. High-contribution samples (red bars), selected via ICon for instruction tuning, exhibit distinct difficulty patterns compared to the full dataset (gray bars). High-contribution ICon samples tend to fall within an appropriate difficulty range.
We further analyzed the characteristics of the high-contribution samples selected by the ICon method. Figure 4 shows the distribution of high-contribution samples in the t-SNE space, where we observe both cross-cluster dispersion and local density. This pattern aligns with the notion in previous studies Liu et al. (2023c); Bukharin and Zhao (2023) that "instruction data should be diverse". However, the distribution is not uniformly scattered. Instead, some clusters exhibit higher density, suggesting that high-contribution samples tend to exhibit certain clustering behavior in the feature space. Figure 5 further compares the distribution of instruction difficulty between high-contribution samples and the full dataset. We use the IFD score Li et al. (2023a, 2024b) as the difficulty metric. The t-test Kendall (1937) result (t = 2.24, p = 0.0251) shows that high-contribution samples are significantly more difficult than the full dataset on average, indicating that the selection method tends to favor more challenging samples. However, as seen in the distribution, high-contribution samples are not exclusively concentrated in the highest difficulty range. Interestingly, while the method favors more difficult samples, some mid-level difficulty samples are retained, and extreme difficulty samples are avoided. This finding challenges the view that more difficult samples are always inherently more beneficial for model training Du et al. (2023); Li et al. (2023a, 2024b, 2024a); Zhang et al. (2025).
7 Conclusion
In this paper, we introduced ICon, a novel method that uses ICL to efficiently assess the contribution of individual training samples without human inductive bias. The method is structured into three stages: constructing an assessment set, computing ICon scores, and training a lightweight selection paradigm for scalable data selection. Extensive experiments on three LLMs across 12 benchmarks and 5 pairwise evaluation sets show that models trained on ICon-selected data outperform those trained on the full dataset, achieving superior performance with fewer samples. ICon also significantly outperforms other widely used data selection methods. We further analyzed the optimal data scale, generalization across datasets, and characteristics of high-contribution samples. As data scale increases, performance improves initially but later declines, indicating an optimal data size. ICon generalizes well across datasets, and its high-contribution samples display diversity while maintaining an appropriate level of difficulty for effective learning.
Limitations
While ICL can be seen as a form of implicit fine-tuning, it captures the impact of individual samples on the model’s behavior, which may not fully replicate the effect of batch training with gradient updates. As a result, some differences in performance may arise. Nevertheless, we believe that ICL provides a useful approximation of sample contribution and can still effectively guide data selection for instruction tuning despite these differences. Furthermore, while our method demonstrates promising results for instruction tuning, we believe it has the potential for broader applicability. For instance, ICon could be adapted to other tasks and domains by modifying the assessment set and the calculation of the ICon score. Exploring these extensions is beyond the scope of this paper and remains a direction for future work.
Appendix A Detailed Evaluation Benchmark for Different Methods
The detailed results for different methods on evaluation benchmarks are shown in Table 5. The methods include Low PPL, Top PPL, Alpagasus, Deita, Superfilter, and ICon. Overall, ICon outperforms existing methods across the evaluated benchmarks.
[…]
IFEval AE Avg GLUE GPQA MMLU TQA Avg ARC BBH HS LQA MuSR WG Avg
Low PPL 21.28 17.42 19.35 56.51 29.80 53.18 38.30 44.45 44.97 40.51 56.32 30.88 37.37 69.14 46.53 41.31
Top PPL 0.97 5.60 3.28 55.29 26.77 44.82 39.98 41.72 43.69 37.01 54.55 23.04 44.10 65.75 44.69 36.80
Alpagasus 9.72 23.07 16.39 57.52 30.81 53.82 38.39 45.14 44.20 41.04 56.13 27.19 37.74 65.43 45.29 40.42
Deita 9.10 21.30 15.20 58.31 27.28 54.47 41.77 45.46 42.66 42.94 55.67 26.57 39.90 67.79 45.92 40.65
Superfilter 13.98 26.53 20.25 55.64 27.78 53.66 41.94 44.76 45.65 41.31 56.56 25.81 37.12 66.22 45.45 41.02
ICon (Ours) 21.16 29.89 25.52 60.58 30.81 52.30 43.16 46.71 44.80 44.15 55.56 28.42 41.64 67.96 47.09 43.37
Table 5: The detailed results of different methods on the evaluation benchmarks. The datasets AE, TQA, HS, LQA, and WG correspond to AlpacaEval, TruthfulQA, HellaSwag, LogicQA, and WinoGrande, respectively. All models are trained on LLaMA3.1-8B with 7,800 selected samples. The best results are highlighted in bold.
Appendix B Detailed Evaluation Benchmark for Different Data Scales
[…]
IFEval AE Avg GLUE GPQA MMLU TQA Avg ARC BBH HS LQA MuSR WG Avg
ICon (1%) 8.31 28.62 18.46 58.47 31.31 55.77 43.11 47.17 49.91 44.51 61.19 23.35 39.37 71.03 48.23 42.91
ICon (5%) 16.21 28.96 22.58 56.91 27.27 53.70 40.51 44.60 47.18 44.23 58.44 24.42 37.75 69.46 46.91 42.09
ICon (10%) 16.08 27.99 22.03 57.86 26.76 53.57 42.63 45.21 45.48 42.81 56.27 26.57 42.01 66.93 46.68 42.08
ICon (15%) 21.16 29.89 25.52 60.58 30.81 52.30 43.16 46.71 44.80 44.15 55.56 28.42 41.64 67.96 47.09 43.37
ICon (20%) 11.95 27.24 19.60 55.07 29.29 53.02 41.71 44.77 43.36 42.08 55.06 26.88 37.36 64.88 44.94 40.66
ICon (25%) 15.63 23.35 19.49 50.84 28.28 51.12 41.26 42.88 42.06 42.13 54.52 28.88 39.77 67.32 45.78 40.43
ICon (30%) 17.72 25.87 21.79 58.57 24.74 52.41 40.66 44.10 41.89 42.32 54.35 26.57 37.63 66.54 44.88 40.77
ICon (50%) 15.00 22.91 18.95 57.75 23.73 51.50 41.12 43.53 41.64 40.33 53.19 28.12 35.27 64.25 43.80 39.57
ICon (75%) 15.17 23.79 19.48 55.11 29.80 46.70 39.94 42.89 39.99 37.24 52.52 28.73 35.12 64.25 42.98 39.03
FULL (100%) 15.68 18.66 17.17 54.48 27.27 43.26 39.60 41.15 40.53 36.46 51.42 26.73 37.51 63.77 42.74 37.95
Table 6: The detailed results of different data scales on the evaluation benchmarks. The datasets AE, TQA, HS, LQA, and WG correspond to AlpacaEval, TruthfulQA, HellaSwag, LogicQA, and WinoGrande, respectively. All models are trained on LLaMA3.1-8B with 7,800 selected samples. The best results are highlighted in bold.
Appendix C Detailed Evaluation Benchmark for Different Dataset
[…]
IFEval AE Avg GLUE GPQA MMLU TQA Avg ARC BBH HS LQA MuSR WG Avg (×$\times 10^{12}$)
ALL 25.43 41.79 33.61 61.13 25.76 48.05 46.91 45.46 39.76 40.56 51.22 27.65 42.41 64.64 44.37 42.94 191.95
[…]
ICon (5%) 26.90 59.52 43.21 62.78 26.26 56.64 47.44 48.28 48.29 44.23 57.73 27.50 41.07 71.35 48.36 47.48 13.06
ICon (10%) 27.67 46.11 36.89 62.89 27.27 55.22 47.21 48.15 43.94 40.57 56.13 29.19 44.18 69.06 47.18 45.79 23.77
ICon (15%) 25.98 44.50 35.24 63.04 25.25 52.63 44.25 46.29 43.52 42.00 54.56 28.11 38.72 67.64 45.76 44.18 30.45
Table 7: The detailed performance of LLaMA3.1-8B trained on WizardLM data in terms of FLOPs and evaluation benchmarks for instruction following, knowledge, and reasoning. The datasets AE, TQA, HS, LQA, and WG correspond to AlpacaEval, TruthfulQA, HellaSwag, LogicQA, and WinoGrande, respectively. The best results of ICon across different training data scales are highlighted in bold.